# Hierarchical Prompting Taxonomy: A Universal Evaluation Framework for Large Language Models Aligned with Human Cognitive Principles

### v4

@@ E.1 HPI for Adaptive HPF
Table 6: HPI (lower is better) of LLMs across datasets (with thresholds) for Adaptive HPF.
@@
Table 7: Performance scores of LLMs across datasetsfor Adaptive HPF.
E.2 Hallucination in Adaptive HPF
@@
For the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations (represented by ’…’). However, by the fourth iteration, Llama-3 8B manages to answer correctly at Level 2. For the CSQA task, prompt-selector exhibits hallucinations initially, shown by Level 4 and Level 0 (not included in HPF) responses. Eventually, it corrects itself by the third iteration, providing the correct answer at Level 2. For the IWSLT task, prompt-selector demonstrates a consistent pattern of hallucinations across multiple iterations. Even though Llama-3 8B attempts the translation at Level 2 multiple times, it ultimately fails to provide a correct translation, indicating a persistent hallucination. For the SamSum task, prompt-selector shows initial hallucinations in the first three iterations (Level 4). However, by the fourth and fifth iterations, the prompt-selector starts producing lower levels. Finally, Llama-3 8B achieves the correct answer at Level 2 in the last iteration .
@@
E.3 Prompt Template for Prompt-Selector
@@
Appendix F Computational Budget
@@
Appendix G Large Language Models Used for Evaluation
@@
Table 8: License information for LLMs used in the experiments.
@@
Appendix H Prompt Templates
H.1 Level 1: Role Prompting
@@
Table 9: Prompt templates of different datasets for Role Prompting.
H.2 Level 2: Zero-shot Chain-of-Thought Prompting
@@
Table 10: Prompt templates of different datasets for Zero-shot Chain-of-Thought Prompting.
H.3 Level 3: Three-Shot Chain-of-Thought Prompting
@@
Table 11: Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting.
H.4 Level 4: Least-to-Most Prompting
@@
Table 12: Prompt templates of different datasets for Least-to-Most Prompting.
H.5 Level 5: Generated Knowledge Prompting
@@
Table 13: Prompt templates of different datasets for Generated Knowledge Prompting.

### v5

@@ B.3. Scoring Method
Table 7. Performance scores of LLMs across datasetsfor Adaptive HPF.
@@
Appendix C Hallucination in Adaptive HPF
@@
For the BoolQ task, the prompt-selector initially struggles, indicated by the iterations where it reaches Level 4 with hallucinations. However, by the fourth iteration, Llama-3 8B manages to answer correctly at Level 2. For the CSQA task, prompt-selector exhibits hallucinations initially, shown by Level 4 and Level 0 (not included in HPF) responses. Eventually, it corrects itself by the third iteration, providing the correct answer at Level 2. For the IWSLT task, prompt-selector demonstrates a consistent pattern of hallucinations across multiple iterations. Even though Llama-3 8B attempts the translation at Level 2 multiple times, it ultimately fails to provide a correct translation, indicating a persistent hallucination. For the SamSum task, prompt-selector shows initial hallucinations in the first three iterations (Level 4). However, by the fourth and fifth iterations, the prompt-selector starts producing lower levels. Finally, Llama-3 8B achieves the correct answer at Level 2 in the last iteration .
@@
C.1. Prompt Template for Prompt-Selector
@@
Appendix D Computational Budget
@@
Appendix E Large Language Models Used for Evaluation
@@
Table 8. License information for LLMs used in the experiments.
@@
Appendix F Prompt Templates
F.1. Level 1: Role Prompting
@@
Table 9. Prompt templates of different datasets for Role Prompting.
@@
F.2. Level 2: Zero-shot Chain-of-Thought Prompting
@@
Table 10. Prompt templates of different datasets for Zero-shot Chain-of-Thought Prompting.
@@
F.3. Level 3: Three-Shot Chain-of-Thought Prompting
@@
Table 11. Prompt templates of different datasets for Three-Shot Chain-of-Thought Prompting.
@@
F.4. Level 4: Least-to-Most Prompting
@@
Table 12. Prompt templates of different datasets for Least-to-Most Prompting.
@@
F.5. Level 5: Generated Knowledge Prompting
@@
Table 13. Prompt templates of different datasets for Generated Knowledge Prompting.